OpenStudy (babyslapmafro):

Please help me solve for x in the following equation. x^3+x^2-2=0 What is the process for solving for x in equations like this and of higher degree? thanks

OpenStudy (anonymous):

Do you know how to factor it?

OpenStudy (agent0smith):

You could check for rational zeroes. Factors of -2, since the coefficient on x^3 is 1.

OpenStudy (agent0smith):

Graphically is another method.

OpenStudy (anonymous):

Typing 'factor(x^3+x^2-2)' at wolframalpha.com coult help as well.

OpenStudy (anonymous):

it would be x^3*x^2=2

OpenStudy (agent0smith):

^that isn't helpful @drewsheesley

OpenStudy (anonymous):

(x-1)*(x^2 +2x +2)=(x-1)*((x+1)^2 +1)=0 therefore x=1 or (x+1)^2=-1 x=1 or x= i-1 or x= -i -1

OpenStudy (anonymous):

put x=1 1+1-2=0 0=0 divide by x-1 ,we get 

OpenStudy (anonymous):

the process is euclidean division by trying out roots so you can factorize and then you will have to learn complex numbers and euler equations, depending on how you want to display the results....